the difference of 4 times a number and 8 is 36?
The difference of 2 times a number and 6 is -24

Answer by checkley71(8403)

4x-8=36
4x=36+8
4x=44
x=44/4
x=11 answer.
---------------------------------
2x-6=-24
2x=-24+6
2x=-18
x=-18/2
x=-9 answer.